This one is really hard for me...can you help me? I have to complete the square, and find the roots of the quadratic equation.
x^2 - 2/3x = 0

Answer by Prithwis(166)

x^2 - 2/3x = 0
=> x^2 - 2*1/3*x + (1/3)^2 = 0 + (1/3)^2
=> (x-1/3)^2 = (1/3)^2
=> (x-1/3) = 1/3 or (x-1/3) = -1/3
=> x = 2/3 or x = 0